OpenStudy (anonymous):

Given sinӨ = 6/11 and sec Ө > 0, find cosӨ and tanӨ

OpenStudy (jtvatsim):

What have you tried to do so far?

OpenStudy (anonymous):

I just don't know what to do at all >_<

OpenStudy (jtvatsim):

OK, so you might want to think of using the trig identity which says: sin^2(x) + cos^2(x) = 1

OpenStudy (jtvatsim):

If you plug in your value of sin(x) = 6/11, you can solve for cos(x).

OpenStudy (anonymous):

Those are the answer options

OpenStudy (campbell_st):

ok... start by drawing a right triangle  find a by using pythagoras' theorem.. then since sec > 0 then cos is.., and since sin > 0 then its a 1st quadrant angle...

OpenStudy (campbell_st):

you can now find cos and tan

OpenStudy (anonymous):

So I can't use sine and cosine to find a? I have to use Pythagoras theorem?

OpenStudy (campbell_st):

if you find a by pythagoras then if \[\sin(x) = \frac{6}{11}\] then \[\cos(x) = \frac{a}{11} ... and... \tan(x) = \frac{6}{a}\]

OpenStudy (anonymous):

Oh I get you

OpenStudy (campbell_st):

its the simplest way to do these questions.. draw the triangle find the missing side check the quadrant find the other ratios...

OpenStudy (anonymous):

So my a would be square root of 85 or 9.22

OpenStudy (campbell_st):

use sqrt 85 since that is one of your answer choices... don't use the decimal. the sqrt 85 will give you an exact value, rather than a decimal approximation.

OpenStudy (anonymous):

So my answer would be a sqrt 85/11 c:

OpenStudy (campbell_st):

yes that will be cos and tan will be \[\tan(x) = \frac{6}{\sqrt{85}}\]

OpenStudy (anonymous):

ohhh okai I see c:

OpenStudy (anonymous):

Thank you n.n

OpenStudy (campbell_st):

so to me cos looks like the 1st option you posted.... not sure of the letter...

OpenStudy (anonymous):

yeah the first one is a cx

OpenStudy (campbell_st):

ps... I used x instead of theta... sorry about that

OpenStudy (anonymous):

it's okai I picked up on that cx